# The KSBA moduli space of stable log Calabi–Yau surfaces

### v2

@@ 3.5. Open Kulikov degenerations
    There exists $10<c_{i}<$ 1 such that $-(K_{Y}+\sum_{i}c_{i}D_{i})$ is big and nef.
@@
The equivalence between (i), (ii), (iii) and (iv) is the equivalence between (1.1), (1.2), (1.3) and (1.4) in [41, Lemma 6.9].
If $W=-(K_{Y}+\sum_{i}c_{i}D_{i})$ is big and nef, then $-K_{Y}=W+\sum_{i}c_{i}D_{i}$ is the sum of a big divisor with an effective divisor, and so $-K_{Y}$ big. This shows that (iv) implies (v). Finally, we show that (v) implies (iii). If $-K_{Y}$ is big, then $\sum_{i}D_{i}=-K_{Y}$ is big, and so the result follows from the proof that (1.1) implies (1.2) in the proof of [41, Lemma 6.9]. ∎
@@
By assumption, the open Kulikov degeneration $\pi:\mathcal{X}\rightarrow\Delta$ is quasi-projective. Hence, by Lemma 3.31, the contraction $f:\mathcal{X}\rightarrow\overline{\mathcal{X}}$ is projective. Thus, by the relative Kodaira Lemma applied to $f:\mathcal{X}\rightarrow\overline{\mathcal{X}}$, there exists an effective divisor $F$ on $\mathcal{X}$ such that $-F$ is $f$-ample. Write $F=a_{i}X^{i}+G$, where $a_{i}$ is an integer and $G$ is a divisor not containing $X^{i}$. As $X^{i}$ is a compact irreducible component of $\mathcal{X}_{0}$, $X^{i}$ is contracted by $f$ to a point and so it follows from the Negativity Lemma [67, Lemma 3.39] that $a_{i}>0$. Restricting $F$ to $X^{i}$, we obtain:
@@
As $K_{\mathcal{X}}=0$, it follows by adjunction that the normal bundle to $X^{i}$ in $\mathcal{X}$ is $K_{X^{i}}$, and so $X^{i}|_{X^{i}}=K_{X^{i}}$. Hence, we have
$-K_{X^{i}}=\frac{1}{a_{i}}(-F|_{X^{i}}+G|_{X^{i}})\,.$ (3.12)
@@
As $X^{i}$ is contracted by $f$ and $-F$ is $f$-ample, it follows that $-F|_{X^{i}}$ is ample on $X^{i}$. On the other hand, as $G$ does not contain $X^{i}$, $G|_{X^{i}}$ is effective. Therefore, $-F|_{X^{i}}+G|_{X^{i}}$ is the sum of an ample divisor with an effective divisor, and so is big. Hence, $-K_{X^{i}}$ is big by the Equation (3.12) above and using the fact that $a_{i}>0$. Finally, we conclude that $(X^{i},\partial X^{i})$ is positive using that (v) implies (i) in Lemma 3.33
@@ 3.6. Mori fans of quasi-projective open Kulikov degenerations
    There is a finite collection of small $\mathbb{Q}$-factorial modifications $f_{i}:V\dashrightarrow V_{i}$ such that each $V_{i}$ satisfies 2)2)2 ) and the Movable cone of relative moving divisors $\mathrm{Mov}(V/W)$ is a union of the cones $f_{i}^{*}(\mathrm{Nef}(V_{i}/W))$.
@@
The total space of a quasi-projective open Kulikov degeneration $\mathcal{X}$ is a Mori dream space over $\overline{\mathcal{X}}$.
@@
This follows from Theorem 3.29, using [23, Corollary 1.3.2] adapted to the formal setup – see [75]. ∎
In what follows we describe the Mori fan associated to an open Kulikov degeneration. For any Mori dream space $V$ over $W$, where $V$ and $W$ are normal algebraic varieties with a projective morphism $f:V\to W$ there is an associated fan in $\mathrm{Pic}(V/W)_{\mathbb{R}}$, known as the Mori fan, with support the cone of $f$-effective divisors, $\mathrm{Eff}(V/W)$ [53, 80]. Using formal Mori theory as in [75], one can similarly construct a Mori fan for a Mori dream space $V$ over $W$, where $V$ and $W$ are formal schemes and $f:V\rightarrow W$ is a projective morphism. The maximal dimensional cones of the Mori fan are in one-to-one correspondence with the birational contractions $g:V\dashrightarrow V^{\prime}$ over $W$ where $V^{\prime}$ is $\mathbb{Q}$-factorial. For such a birational contraction, the corresponding maximal cone of the Mori fan is given by $g^{*}\mathrm{Nef}(V^{\prime}/W)+\mathrm{Ex}(g)$, where $\mathrm{Ex}(g)$ is the cone in $\mathrm{Pic}(V/W)_{\mathbb{R}}$ spanned by all $g$-exceptional divisors. For a birational projective morphism $f:V\to W$, we have $\mathrm{Eff}(V/W)=\mathrm{Pic}(V/W)_{\mathbb{R}}$, and so the Mori fan is a complete fan in $\mathrm{Pic}(V/W)_{\mathbb{R}}$ in this case. Given an open Kulikov degeneration, with affinization $f\colon\mathcal{X}\to\overline{\mathcal{X}}$, $\mathcal{X}$ is a Mori dream space over $\overline{\mathcal{X}}$ by Proposition 3.36, and so we have an associated complete Mori fan in $\mathrm{Pic}(\mathcal{X}/\overline{\mathcal{X}})$, denoted by $\mathrm{MF}(\mathcal{X}/\overline{\mathcal{X}})$ in the following sections.

### v3

@@ 3.4. Open Kulikov degenerations
    There exists $0<c_{i}<1$ such that $(K_{Y}+\sum_{i}c_{i}D_{i})$ is big and nef.
@@
The equivalence between i), ii), iii) and iv) is the equivalence between (1.1), (1.2), (1.3) and (1.4) in [44, Lemma 6.9].
If $W=-(K_{Y}+\sum_{i}c_{i}D_{i})$ is big and nef, then $K_{Y}=W+\sum_{i}c_{i}D_{i}$ is the sum of a big divisor with an effective divisor, and so $K_{Y}$ big. This shows that iv) implies v). Finally, we show that v) implies iii). If $K_{Y}$ is big, then $\sum_{i}D_{i}=-K_{Y}$ is big, and so the result follows from the proof that (1.1) implies (1.2) in the proof of [44, Lemma 6.9]. ∎
@@
By assumption, the open Kulikov degeneration $\pi:\mathcal{X}\rightarrow\Delta$ is quasi-projective. Hence, by Lemma 3.24, the contraction $f:\mathcal{X}\rightarrow\mathcal{X}^{\mathrm{can}}$ is projective. Thus, by the relative Kodaira Lemma applied to $f:\mathcal{X}\rightarrow\mathcal{X}^{\mathrm{can}}$, there exists an effective divisor $F$ on $\mathcal{X}$ such that F-F is $f$-ample. Write $F=a_{i}X^{i}+G$, where $a_{i}$ is an integer and $G$ is a divisor not containing $X^{i}$. Since $X^{i}$ is a compact irreducible component of $\mathcal{X}_{0}$, $X^{i}$ is contracted by $f$ to a point and so it follows from the Negativity Lemma [73, Lemma 3.39] that $a_{i}>0$. Restricting $F$ to $X^{i}$, we obtain:
@@
Since $K_{\mathcal{X}}=0$, it follows by adjunction that the normal bundle to $X^{i}$ in $\mathcal{X}$ is $K_{X^{i}}$, and so $X^{i}|_{X^{i}}=K_{X^{i}}$. Hence, we have
$K_{X^{i}}=\frac{1}{a_{i}}(-F|_{X^{i}}+G|_{X^{i}})\,.$ (3.8)
@@
As $X^{i}$ is contracted by $f$ and F-F is $f$-ample, it follows that $F|_{X^{i}}$ is ample on $X^{i}$. On the other hand, as $G$ does not contain $X^{i}$, $G|_{X^{i}}$ is effective. Therefore, $F|_{X^{i}}+G|_{X^{i}}$ is the sum of an ample divisor with an effective divisor, and so is big. Hence, $K_{X^{i}}$ is big by the Equation (3.8) above and using the fact that $a_{i}>0$. Finally, we conclude that $(X^{i},\partial X^{i})$ is positive using that v) implies i) in Lemma 3.26
3.5. Generic open Kulikov degenerations
In this section, we introduce a notion of generic open Kulikov surface, and prove in Theorem 3.31 that every open Kulikov surface admits a generic $d$-semistable locally trivial deformation.
Let $\mathcal{X}_{0}$ be an open Kulikov surface, with dual intersection complex $(P,\mathscr{P})$. Let $\hat{\mathcal{X}}_{0}$ be a compactification of $\mathcal{X}_{0}$ defined by choosing a log Calabi–Yau compactification of every non-compact irreducible component of $\mathcal{X}_{0}$. The boundary $\partial{\hat{\mathcal{X}}}_{0}:=\hat{\mathcal{X}}_{0}\setminus\mathcal{X}_{0}$ is a cycle of rational curves. Let $\widetilde{\Lambda}_{\mathcal{X}_{0}}\subset\mathrm{Pic}(\hat{\mathcal{X}}_{0})$ be the subgroup of line bundles $L$ on $\hat{\mathcal{X}}_{0}$ such that $\mathrm{deg}(L|_{C})=0$ for every irreducible component $C$ of $\partial\hat{\mathcal{X}}_{0}$. Different compactifications $\hat{\mathcal{X}}_{0}$ differ by corner blow-ups, and so the corresponding lattices $\widetilde{\Lambda}_{\mathcal{X}_{0}}$ can be canonically identified, and so the lattice $\widetilde{\Lambda}_{\mathcal{X}_{0}}$ is canonically associated to $\mathcal{X}_{0}$. Moreover, it follows from Lemma 3.10 that if $\mathcal{X}_{0}^{\prime}$ is a locally trivial deformation of $\mathcal{X}_{0}$, then we have a canonical identification $\widetilde{\Lambda}_{\mathcal{X}_{0}^{\prime}}\simeq\widetilde{\Lambda}_{\mathcal{X}_{0}}$. We fix an orientation of $\partial\mathscr{P}$, which induces an isomorphism $\mathrm{Pic}^{0}(\partial\hat{\mathcal{X}}_{0})\simeq\mathbb{C}^{\star}$, where $\mathrm{Pic}^{0}(\partial\hat{\mathcal{X}}_{0})$ is the group of line bundles on $\partial\hat{\mathcal{X}}_{0}$, which are of degree zero in restriction to every irreducible component of $\partial\hat{\mathcal{X}}_{0}$. Every locally trivial deformation $\mathcal{X}_{0}^{\prime}$ of $\mathcal{X}_{0}$ defines a period point $\psi_{\mathcal{X}_{0}^{\prime}}\in\mathrm{Hom}(\widetilde{\Lambda}_{\mathcal{X}_{0}},\mathbb{C}^{\star})$, given by
$\psi_{\mathcal{X}_{0}^{\prime}}(L)=[L|_{\partial\hat{\mathcal{X}_{0}}}]\in\mathrm{Pic}^{0}(\partial{\hat{\mathcal{X}}}_{0})\simeq\mathbb{C}^{\star}$
for every $L\in\widetilde{\Lambda}_{\mathcal{X}_{0}}$. For every interior vertex $v$ of $\mathscr{P}$, corresponding to a compact irreducible component $X^{v}$ of $\mathcal{X}_{0}$, define $\xi_{v}=\sum_{v^{\prime}\in\mathscr{P}^{[0]}}(D_{vv^{\prime}}-D_{v^{\prime}v})\in\widetilde{\Lambda}_{\mathcal{X}_{0}}$, where $D_{vv^{\prime}}=X^{v}\cap X^{v^{\prime}}$. For example, when $\mathcal{X}_{0}$ admits a smoothing $\mathcal{X}$, we have $\xi_{v}=[\mathcal{O}_{\mathcal{X}}(-X^{v})|_{\mathcal{X}_{0}}]$.
Lemma 3.28.
A locally trivial deformation $\mathcal{X}_{0}^{\prime}$ of $\mathcal{X}_{0}$ is d-semistable if and only if $\psi_{\mathcal{X}_{0}^{\prime}}(\xi_{v})=1$ for every interior vertex $v$ of $\mathscr{P}$.
Proof.
The normal crossing surface $\mathcal{X}_{0}^{\prime}$ is $d$-semistable if $T^{1}_{\mathcal{X}^{\prime}_{0}}=\mathcal{E}xt^{1}(\Omega^{1}_{\mathcal{X}_{0,\psi}},\mathcal{O}_{\mathcal{X}_{0,\psi}})\simeq\mathcal{O}_{\mathcal{X}^{\prime}_{0,\mathrm{sing}}}$, where $\mathcal{X}^{\prime}_{0,\mathrm{sing}}$ is the double locus of $\mathcal{X}^{\prime}_{0}$, see [35, 37]. The sheaf $T^{1}_{\mathcal{X}^{\prime}_{0}}$ is an invertible sheaf on $\mathcal{X}^{\prime}_{0,\mathrm{sing}}$ and its class lies in $\operatorname{Pic}^{0}(\mathcal{X}^{\prime}_{0,\mathrm{sing}})=\mathrm{Hom}(H_{1}(\Gamma(\mathcal{X}^{\prime}_{0,\mathrm{sing}}),\mathbb{C}^{*}))\simeq(\mathbb{C}^{*})^{g}$, where $\Gamma(\mathcal{X}^{\prime}_{0,\mathrm{sing}})$ is the dual graph of $\mathcal{X}^{\prime}_{0,\mathrm{sing}}$ and $g$ is its genus. The group $H_{1}(\Gamma(\mathcal{X}^{\prime}_{0,\mathrm{sing}}))$ is isomorphic to $\mathbb{Z}^{g}$, with a basis consisting of the loops $\Gamma(\partial X^{v})$ corresponding to the boundaries of the compact components $X^{v}$. As in [7, §4A], the restriction $T^{1}_{\mathcal{X}^{\prime}_{0,\psi}}|_{\partial X^{v}}$ is $\psi_{\mathcal{X}_{0}^{\prime}}(\xi_{v})\in\operatorname{Pic}^{0}(\partial X^{v})=\mathbb{C}^{*}$. It follows that $T^{1}_{\mathcal{X}^{\prime}_{0}}=1$ in $\mathrm{Pic}^{0}(\mathcal{O}_{\mathcal{X}_{0}^{\prime},\mathrm{sing}})$ if and only if all $\psi_{\mathcal{X}_{0}^{\prime}}(\xi_{v})=1$. ∎
Denote by $\Xi$ the saturation in $\widetilde{\Lambda}_{\mathcal{X}_{0}}$ of the lattice spanned by the classes $\xi_{v}$ for all interior vertices $v$ of $\mathscr{P}$. Let $\Lambda_{\mathcal{X}_{0}}:=\widetilde{\Lambda}_{\mathcal{X}_{0}}/\Xi$, so that by Lemma 3.28, the torus $\mathrm{Hom}(\Lambda_{\mathcal{X}_{0}},\mathbb{C}^{\star})$ contains period points of $d$-semistable locally trivial deformations of $\mathcal{X}_{0}$.
Lemma 3.29.
There exists a family $\widetilde{\mathcal{X}}_{0}\rightarrow\mathrm{Hom}(\Lambda_{\mathcal{X}_{0}},\mathbb{C}^{\star})$ of $d$-semistable locally trivial deformations of $\mathcal{X}_{0}$, such that the induced period map $\mathrm{Hom}(\Lambda_{\mathcal{X}_{0}},\mathbb{C}^{\star})\rightarrow\mathrm{Hom}(\Lambda_{\mathcal{X}_{0}},\mathbb{C}^{\star})$ is the identity map.
Proof.
Choosing a toric model for each irreducible component of $\mathcal{X}_{0}$, we can define as in [7, §7B] a corresponding glueing complex $\mathcal{G}:\mathbb{Z}^{E+Q}\rightarrow(\mathbb{Z}^{2})^{V}$, where $E$ is the number of edges of $\mathscr{P}$ and $V$ is the number of faces of $\mathscr{P}$, and a universal family of locally trivial deformations of $\mathcal{X}_{0}$ over the torus $\mathrm{Hom}(H^{0}(\mathcal{G}),\mathbb{C}^{\star})$. Arguing as in the proof of [7, Theorem 7.9], we have an isomorphism
$\widetilde{\Lambda}_{\mathcal{X}_{0}}$ $\xrightarrow{\quad\sim\quad}H^{0}(\mathcal{G})$
$L$ $\longmapsto((L\cdot C_{e})_{e},(L\cdot E_{i})_{i})\,$,
where $C_{e}$ are the irreducible components of the double locus of $\hat{\mathcal{X}}_{0}$ corresponding to the edges $e$ of $\mathscr{P}$, and $E_{i}$ are the exceptional curves of the toric models. Moreover, as in the proof of [7, Proposition 7.11], the induced isomorphism of tori $\mathrm{Hom}(H^{0}(\mathcal{G}),\mathbb{C}^{\star})\simeq\mathrm{Hom}(\widetilde{\Lambda}_{\mathcal{X}_{0}},\mathbb{C}^{\star})$ agrees with the period map of the universal family. We obtain $\widetilde{\mathcal{X}}_{0}\rightarrow T$ by restriction of the universal family to the subtorus $\mathrm{Hom}(\Lambda_{\mathcal{X}_{0}},\mathbb{C}^{\star})$ induced by the surjection $\widetilde{\Lambda}_{\mathcal{X}_{0}}\rightarrow\Lambda_{\mathcal{X}_{0}}=\widetilde{\Lambda}_{\mathcal{X}_{0}}/\Xi$. ∎
Definition 3.30.
A $d$-semistable open Kulikov surface $\mathcal{X}_{0}$ is called generic if $\psi_{\mathcal{X}_{0}}(\gamma)\neq 1$ for every $\gamma\in\Lambda_{\mathcal{X}_{0}}\setminus\{0\}$. An open Kulikov degeneration $\mathcal{X}\rightarrow\Delta$ is called generic if its central fiber $\mathcal{X}_{0}$ is a generic $d$-semistable open Kulikov surface.
Theorem 3.31.
Every open Kulikov surface admits a generic $d$-semistable locally trivial deformation.
Proof.
This follows from the existence of a family of $d$-semistable locally trivial deformations of $\mathcal{X}_{0}$ as in Lemma 3.29. ∎
Lemma 3.32.
Let $\mathcal{X}_{0}$ be a generic $d$-semistable open Kulikov surface. Then, no irreducible component of $\mathcal{X}_{0}$ contains an internal (2)(-2)-curve.
Proof.
For every vertex $v$ of $\mathscr{P}$, denote by $\Lambda_{X^{v}}\subset\mathrm{Pic}(X^{v})$ the subgroup of line bundles $L$ on $X^{v}$ such that $\mathrm{deg}(L|_{C})=0$ for every irreducible component $C$ of $\partial X^{v}$. The period point $\psi_{X^{v}}\in\mathrm{Hom}(\Lambda_{X^{v}},\mathbb{C}^{\star})$ of $X^{v}$ is defined by $\psi_{X^{v}}(L)=[L|_{\partial X^{v}}]$, as in [45].
By Lemma 3.10, every line bundle $L\in\Lambda_{X^{v}}$ uniquely extends to a line bundle $\widetilde{L}\in\widetilde{\Lambda}_{\mathcal{X}_{0}}$, which is trivial in restriction to every irreducible component $X^{v^{\prime}}$ with $v^{\prime}\neq v$. Assume that $X^{v}$ contains an internal (2)(-2)-curve $E$. Then, the line bundle $\mathcal{O}_{X^{v}}(E)$ uniquely extends to a line bundle $\mathcal{O}_{\mathcal{X}_{0}}(E)$ on $\mathcal{X}_{0}$, which is trivial in restriction to every irreducible component $X^{v^{\prime}}$ with $v^{\prime}\neq v$. For every interior vertex $v^{\prime}$, we have $\xi_{v^{\prime}}\cdot E=0$, whereas $\mathcal{O}_{\mathcal{X}_{0}}(E)\cdot E=-2$. So, we have $\mathcal{O}_{\mathcal{X}_{0}}(E)\notin\Xi$, that is, $\mathcal{O}_{\mathcal{X}_{0}}(E)$ is non-zero in $\Lambda_{\mathcal{X}_{0}}=\widetilde{\Lambda}_{\mathcal{X}_{0}}/\Xi$. We have $\psi_{X^{v}}(\mathcal{O}_{X^{v}}(E))=\psi_{\mathcal{X}_{0}}(\mathcal{O}_{\mathcal{X}_{0}}(E))$, but $\psi_{X^{v}}(\mathcal{O}_{X^{v}}(E))=1$, so $\psi_{\mathcal{X}_{0}}(\mathcal{O}_{\mathcal{X}_{0}}(E))=1$, in contradiction with the assumption that $\mathcal{X}_{0}$ generic. ∎
Lemma 3.33.
Let $\mathcal{X}\rightarrow\Delta$ be a generic open Kulikov degeneration. Then, the following holds:
  • i)
    The generic rigid-analytic fiber $\mathcal{X}_{\eta}$ of $\mathcal{X}$ does not contain any (2)(-2)-curve.
  • ii)
    The affinization induces an isomorphism $\mathcal{X}_{\eta}\simeq\mathcal{X}_{\eta}^{\mathrm{can}}$.
  • iii)
    $\mathcal{X}^{\mathrm{can}}\rightarrow\Delta$ is a formal smoothing of $\mathcal{X}_{0}^{\mathrm{can}}$.
Proof.
Assume by contradiction that $\mathcal{X}_{\eta}$ contains a (2)(-2)-curve $E_{\eta}$. Let $E$ denote the closure of $E_{\eta}$ in $\mathcal{X}$. Since $\mathcal{X}$ is regular, $E$ is a Cartier divisor on $\mathcal{X}$. Moreover, $E$ does not contain any irreducible component of $\mathcal{X}_{0}$, and so is flat over $\Delta$. Thus, $E$ is a relative effective Cartier divisor as in [93, Tag 056P]. Hence, by [93, Tag 056Q], the pullback $E_{0}$ of $E$ to $\mathcal{X}_{0}$ is a Cartier divisor in $\mathcal{X}_{0}$, corresponding to the line bundle $\mathcal{O}_{\mathcal{X}_{0}}(E_{0})=\mathcal{O}_{\mathcal{X}}(E)|_{\mathcal{X}_{0}}$. By construction, we have $\mathcal{O}_{\mathcal{X}_{0}}(E_{0})\in\Lambda_{\mathcal{X}_{0}}$, $\psi_{\mathcal{X}_{0}}(\mathcal{O}_{\mathcal{X}_{0}}(E_{0}))=1$ and $\mathcal{O}_{\mathcal{X}_{0}}(E_{0})\cdot E_{0}=-2$. Moreover, for every interior vertex $v$, we have $\mathcal{O}_{\mathcal{X}}(X^{v})\cdot E_{\eta}=0$, and so $\mathcal{O}_{\mathcal{X}_{0}}(X^{v})\cdot E_{0}=0$. It follows that $\mathcal{O}_{\mathcal{X}_{0}}(E_{0})\notin\Xi$, that is, $\mathcal{O}_{\mathcal{X}_{0}}(E_{0})$ is non-zero in $\Lambda_{\mathcal{X}_{0}}=\widetilde{\Lambda}_{\mathcal{X}_{0}}/\Xi$, in contradiction with the assumption that $\mathcal{X}_{0}$ is generic. This proves Lemma 3.33 i). Items ii)-iii) then follow from i) by Theorem 3.22. ∎
Remark 3.34.
Let $\mathcal{X}\rightarrow\Delta$ be a generic open Kulikov degeneration. By Lemma 3.33, $\mathcal{X}^{\mathrm{can}}$ is an affine formal smoothing, and so it follows from Elkik’s theorem [30, Theorem 7, p582] – see also [96, Proposition 3.3.1], that $\mathcal{X}^{\mathrm{can}}$ is algebraizable over $\mathrm{Spec}\,\mathbb{C}[\![t]\!]$. Moreover, $\mathcal{X}\rightarrow\mathcal{X}^{\mathrm{can}}$ is a formal blow-up, and so is algebraizable in the category of algebraic spaces over $\mathrm{Spec}\,\mathbb{C}[\![t]\!]$ by [19, Theorem 3.2]. Since $H^{2}(\mathcal{X}_{0},\mathcal{O}_{\mathcal{X}_{0}})=0$, and $\mathcal{X}\rightarrow\mathcal{X}^{\mathrm{can}}$ is projective, it follows that $\mathcal{X}\rightarrow\mathcal{X}^{\mathrm{can}}$ is in fact algebraizable in the category of schemes over $\mathrm{Spec}\,\mathbb{C}[\![t]\!]$.
@@ 3.6. Mori fans of quasi-projective open Kulikov degenerations
    There is a finite collection of small $\mathbb{Q}$-factorial modifications $f_{i}:V\dashrightarrow V_{i}$ such that each $V_{i}$ satisfies ii) and the Movable cone of relative moving divisors $\mathrm{Mov}(V/W)$ is a union of the cones $f_{i}^{*}(\mathrm{Nef}(V_{i}/W))$.
@@
The total space of a quasi-projective open Kulikov degeneration $\mathcal{X}$ is a Mori dream space over $\mathcal{X}^{\mathrm{can}}$.
@@
This follows from Theorem 3.22, using [24, Corollary 1.3.2] adapted to the formal setup – see [80]. ∎
In what follows, we describe the Mori fan associated with an open Kulikov degeneration. For any Mori dream space $V$ over $W$, where $V$ and $W$ are normal algebraic varieties with a projective morphism $f:V\to W$, there is an associated fan in $\mathrm{Pic}(V/W)_{\mathbb{R}}$, known as the Mori fan. Its support is given by the cone of $f$-effective divisors, $\mathrm{Eff}(V/W)$ [58, 85]. Using formal Mori theory, as developed in [80], one can similarly construct a Mori fan for a Mori dream space $V$ over $W$, where $V$ and $W$ are formal schemes and $f:V\rightarrow W$ is a projective morphism.
The maximal dimensional cones of the Mori fan are in one-to-one correspondence with the birational contractions $g:V\dashrightarrow V^{\prime}$ over $W$ where $V^{\prime}$ is $\mathbb{Q}$-factorial. For such a birational contraction, the corresponding maximal cone of the Mori fan is given by $g^{*}\mathrm{Nef}(V^{\prime}/W)+\mathrm{Ex}(g)$, where $\mathrm{Ex}(g)$ is the cone in $\mathrm{Pic}(V/W)_{\mathbb{R}}$ spanned by all $g$-exceptional divisors. For a birational projective morphism $f:V\to W$, we have $\mathrm{Eff}(V/W)=\mathrm{Pic}(V/W)_{\mathbb{R}}$, and so the Mori fan is a complete fan in $\mathrm{Pic}(V/W)_{\mathbb{R}}$ in this case. Given an open Kulikov degeneration, with affinization $f\colon\mathcal{X}\to\mathcal{X}^{\mathrm{can}}$, the space $\mathcal{X}$ is a Mori dream space over $\mathcal{X}^{\mathrm{can}}$ by Proposition 3.36. Consequently, we have an associated complete Mori fan in $\mathrm{Pic}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$, denoted by $\mathrm{MF}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ in the following sections. Additionally, we note that we have $\mathrm{Pic}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})\simeq\mathrm{Pic}(\mathcal{X}_{0})$ by Lemma 3.23.